OpenStudy (anonymous):

It takes a boy 40 minutes to mow the lawn, but his sister can mow it in 60 minutes. How long would it take them to mow the lawn if they worked together, using two lawn mowers?

OpenStudy (anonymous):

24 minutes.. is the answer correct?

hero (hero):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Sorry.. boy will mow 1/40 lawn in 1 min.. girl will mow 1/60 lawn in 1 min.. so togther, 1/40 + 1/60 = 1/24 lawn would be mowed in 1 min. So 24 minutes...

OpenStudy (anonymous):

it is thanks! How did you go about doing it? I'm confused

OpenStudy (anonymous):

Perfect Thanks!!

hero (hero):

Alternative method: B = Time it takes brother to mow S = Time it takes sister to mow t = time it takes both: \[\large\frac{B \times S}{B + S} = t\]